OpenStudy (anonymous):

Simplify. x/4x+x^2????? :)

OpenStudy (anonymous):

PLEASE! SOMEONE!!!! I NEED HELP!

OpenStudy (callisto):

\[\frac{x}{4x+x^2} = \frac{x}{x(4+x)}\]Cancel the common factor of numerator and denominator, what do you get?

OpenStudy (anonymous):

The common factor would be x, right?

OpenStudy (callisto):

Yes~

OpenStudy (anonymous):

So... It would be 1/x+4?

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (callisto):

welcome :)